OpenStudy (beccab003):

Radical Multiplication Problem Help! Medal to best answer. Problem will be posted below.

OpenStudy (beccab003):

\[2+\sqrt{6}\] OVER \[2-\sqrt{6}\]

OpenStudy (danjs):

hi

OpenStudy (danjs):

\[\frac{ 2 + \sqrt{6} }{ 2 - \sqrt{6} }\]

OpenStudy (danjs):

For these, you want to clear a radical from the denominator for a formal answer... Do that by multiplying both the top and bottom by 2 + root(6) like this...

OpenStudy (beccab003):

HI!! Exactly. I just didn't know hot to make that line.

OpenStudy (danjs):

\[\frac{ 2+\sqrt{6} }{ 2-\sqrt{6} }*\frac{2+\sqrt{6} }{ 2+\sqrt{6} } = \]

OpenStudy (danjs):

notice you are not changing the original fraction, you are multiplying it by 1,

OpenStudy (danjs):

2 plus root 6 over 2 plus root 6 is the same as 1

OpenStudy (beccab003):

Do you always use the numerator?

OpenStudy (danjs):

you use the opposite sign in the denominator. there is a -, use a plus for this one.. watch what happens to the denominator when you distribute them together.. foil

OpenStudy (danjs):

\[\frac{ 2+\sqrt{6} }{ 2-\sqrt{6} }*\frac{2+\sqrt{6} }{ 2+\sqrt{6} } = \frac{ 4 + 2\sqrt{6}+2\sqrt{6}+\sqrt{6}\sqrt{6} }{ 4 +2\sqrt{6}-2\sqrt{6} -\sqrt{6}\sqrt{6} }\]

OpenStudy (beccab003):

so I do foil like this: \[(2+\sqrt{6})(2-\sqrt{6}\]

OpenStudy (danjs):

yeah, notice the denominator now has +2root6 - 2 root 6,, the root dissapears

OpenStudy (danjs):

the general trick is , if you have \[a - \sqrt{b}\] in the denominator, multiply the top and bottom by the opposite sign a + root(b)

OpenStudy (beccab003):

okay. so doing foil.... \[4-\sqrt{36}\]

OpenStudy (danjs):

yep, but notice the root simplifies, \[\sqrt{n}\sqrt{n} = n\]

OpenStudy (beccab003):

So...how does it look different?

OpenStudy (danjs):

\[\frac{ 2+\sqrt{6} }{ 2-\sqrt{6} }*\frac{2+\sqrt{6} }{ 2+\sqrt{6} } = \frac{ 4 + 2\sqrt{6}+2\sqrt{6}+\sqrt{6}\sqrt{6} }{ 4 +2\sqrt{6}-2\sqrt{6} -\sqrt{6}\sqrt{6} } = \frac{ 4 + 4\sqrt{6}+6 }{ 4 - 6 }\]

OpenStudy (danjs):

\[\frac{ 10 + 4\sqrt{6} }{ -2 }\]

OpenStudy (danjs):

now factor a -2 from the top

OpenStudy (beccab003):

I don't really understand what the problem is doing from your third up message.

OpenStudy (danjs):

\[\frac{ 2(5+2\sqrt{6}) }{ -2 } = \frac{ -5 - 2\sqrt{6} }{ 1 }\]

OpenStudy (danjs):

k, ill go back

OpenStudy (danjs):

\[\frac{ 2+\sqrt{6} }{ 2-\sqrt{6} }*\frac{2+\sqrt{6} }{ 2+\sqrt{6} } = \]

OpenStudy (beccab003):

I just don't really understand what happens form using foil to the fraction stuff. I get the first part you sent.

OpenStudy (beccab003):

*from

OpenStudy (danjs):

when you foil distribute the top you get \[4 + 2\sqrt{6 }+ 2\sqrt{6 } + \sqrt{6}\sqrt{6}\] you see that part?

OpenStudy (danjs):